OpenStudy (anonymous):

evaluate the expression (4x^3)^2 for x=2

OpenStudy (anonymous):

\[(4x^3)^2, x=2\]

mathslover (mathslover):

\[\huge{(4x^3)^2}\] \[\huge{[4(2^3)]^2}\] \[\huge{(4*8)^2}\]

mathslover (mathslover):

can you do the left steps ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[(4x^{3})^{2}\]Use the rule: \[(x^{a})^{b} = a^{a \times b}\]\[(4x^{3})^{2} = 4^{2}x^{3 \times 2}\]

OpenStudy (anonymous):

Now can you simplify \(\huge4^{2}x^{3 \times 2}\space?\)

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

np :)